OpenStudy (anonymous):

Joselyn is a manager at a sign painting company. She has three painters, Allen, Brianne, and Charles. Allen can complete a large project in 16 hours. Brianne can complete the same sized project in 18 hours. Charles is new, so no one knows how long it will take him. Joselyn assigns them all a large project to complete together. Explain to Joselyn how this project can tell her how long it would take Charles if he worked by himself. Use complete sentences.

OpenStudy (anonymous):

@Reaper534 @jdoe0001 @RadEn @Jordan7 @superdavesuper Help me please

OpenStudy (judygreeneyes):

Haha -- I have a BA in Math and still hate these kinds of questions and the ones about trains going opposite directions!! This is the source I always go back to: http://www.purplemath.com/modules/workprob.htm Allen can complete 1/16 of the project per hour Brianne can complete 1/18 of the project per hour Charles can complete 1/x Together they can accomplish 1/16 + 1/18 + 1/X per hour. we need a common denominator, which is 288X =\[\frac{ 18x + 16x + 288 }{ 288x }\] Choose t as the time it takes them all, so 1/t is the hourly rate. 1/t = \[\frac{ 18x + 16x + 288 }{ 288x }\] or t = \[\frac{ 288x }{ 18x + 16x +288}\] Once the project is completed, Jocelyn will have t, so she will only need to solve for x in this equation to find out the rate 1/x for Charles working by himself (assuming he gets good training!)

OpenStudy (anonymous):

@judygreeneyes Thank you so much! You're a life saver :D

OpenStudy (judygreeneyes):

You're very welcome!!

OpenStudy (anonymous):

what does the final answer come out to be then?

OpenStudy (anonymous):

What is the final answer?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

^^^^!!!!!!!!